OpenStudy (anonymous):

cos ( pi/4 + sin^-1 sqrtx/2 ) in terms of x

OpenStudy (anonymous):

\[\cos ( \frac{ \pi }{ 4 } =\sin^{-1} \frac{ \sqrt{x} }{ 2 } ) \]

OpenStudy (anonymous):

something like that , the question

OpenStudy (anonymous):

it should be + symbol not =

hartnn (hartnn):

\(\large \cos(A+B)=\cos A\cos B-\sin A\sin B\) so what will be \(\cos(\pi/4 +y)=....?\) where, A = pi/4 , B = y = \(\sin^{-1}(\sqrt x/2)\)

OpenStudy (anonymous):

but , later on the cos will get abt 0.7071

hartnn (hartnn):

cos pi/4 = sin pi/4 = 1/sqrt 2 yes.

OpenStudy (anonymous):

if cos 45 = 0.7071

OpenStudy (anonymous):

okay , i get it now

OpenStudy (anonymous):

so the answer will be \[\frac{ \sqrt{(4-x)} }{ 2.\sqrt{2} } - \frac{ \sqrt{x} }{ 2.\sqrt{2} }\]

OpenStudy (anonymous):

right ?

hartnn (hartnn):

how?

OpenStudy (anonymous):

err ,from the \[\sin^{-1} \frac{ \sqrt{x} }{ 2 }\] , i find the triangle

hartnn (hartnn):

ok, yes! you're correct :)

hartnn (hartnn):

just checking whether you actually solved it :)

OpenStudy (anonymous):



OpenStudy (anonymous):

ok ok thank you

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

erm , here's another question , i just want to make sure what i'm doing is right

OpenStudy (anonymous):

\[\sec ( \cot ^ {-1} 2x - \frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

should we use the cofunction identities ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes! sec (A-pi/2) = csc A

OpenStudy (anonymous):

but the question want in terms of x , still need to use that formula ?

hartnn (hartnn):

yes, so your Q becomes csc(cot inverse 2x)

OpenStudy (anonymous):

okay , i just know until that , after that what should we do ?

hartnn (hartnn):

cot inverse 2x = y so you need csc y make a triangle...

hartnn (hartnn):

cot y = 2x

OpenStudy (anonymous):



OpenStudy (anonymous):

this is the triangle

hartnn (hartnn):

absolutely correct just find csc from that

OpenStudy (anonymous):

how abt the pi ? should we just leave it like that ?

hartnn (hartnn):

we already dealt with pi ? sec (cot inv 2x - pi/2) = csc (cot inv 2x)

OpenStudy (anonymous):

oh okay , i forgot abt that heheh

OpenStudy (anonymous):

lo the last answer is \[\sqrt{(1+4x^{2})}\]

OpenStudy (anonymous):

i mean so

hartnn (hartnn):

yes! correct :)

OpenStudy (anonymous):

okay , thank you again , if i need you i'll message you :)

hartnn (hartnn):

sure :)